Grade 12th passAlgebra

Find the area of the following circles, given that:
Radius = 14 mm (Take π = 22/7)

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Given, radius of circle = 14 mm
Then,
Area of the circle =πr2
= 22/7 × 142
= 22/7 × 196
= 22 × 28
= 616 mm2

Thanks